# Explicit monotone stable super-time-stepping methods for finite time singularities

### v1

@@ 5.2.1 Combinatorics Formulation
when applied to problem 19 if and only if
@@
Plugging into the RKL1 method 11, we have that
@@
where $[M^{k}u^{t}]_{i}$ denotes the $i$-th element of the vector $M^{k}u^{t}$. Using induction, it can be shown that
$[M^{k}u^{t}]_{i}=\frac{1}{dx^{2k}}\sum_{j=0}^{2k}\binom{2k}{j}(-1)^{j}u_{i+j-k}^{t}=\frac{1}{dx^{2k}}\sum_{j=-k}^{k}\binom{2k}{j+k}(-1)^{j+k}u_{i+j}^{t}.$
@@ 5.2.2 Estimation of the Alternating Sum and Conclusion
In particular, it gives us a form of generalized hypergeometric functions that is always non-negative. Our goal is to write the alternating sum 20 as a generalized hypergeometric function, with the hope that the coefficients are compatible with Clausen’s formula and directly apply the theorem to conclude. Fortunately, this process indeed gives us the desired result using the calculations in the following lemma.
@@
(a) The $s$-stage RKL1 method 11 applied to constant coefficient heat equation under periodic or homogeneous Neumann boundary condition is monotone if and only if
@@
(b) The $s$-stage RKL2 method 12 applied to constant coefficient heat equation under periodic or homogeneous Neumann boundary condition is monotone if and only if
@@
(a) The $s$-stage RKL1 method 11 applied to constant coefficient semilinear heat equationwith Strang splitting is monotone if and only if it satisfies the stability constraint
@@
(b) The $s$-stage RKL2 method 12 applied to constant coefficient semilinear heat equation with Strang splitting is monotone if and only if it satisfies the stability constraint
@@ 5.3 Runge–Kutta–Gegenbauer Methods Monotonicity
The $s$-stage RKG1 method 13 applied to constant coefficient heat equation under periodic or homogeneous Neumann boundary condition is monotone if it satisfies its stability constraint
@@
Also notice that $C_{s}^{(1/2)}(z)=P_{s}(z)$ is exactly the $s$-th degree Legendre polynomial. The $s$-stage RKL1 monotonicity is equivalent to the statement that $\Big{[}P_{s}\Big{(}I+x\tilde{M}\Big{)}u(t)\Big{]}_{i}$ is a non-negative finite linear combination of $[u]_{j}$’s, where $[\tilde{M}u]_{i}:=[u]_{i+1}-2[u]_{i}+[u]_{i-1}$, under the condition $0\leq x\leq\frac{1}{2}$.
@@
$C_{s}^{(\lambda)}(I+x\tilde{M})=(I+x\tilde{M})C_{s-1}^{(\lambda)}(I+x\tilde{M})+(s+1)C_{s}^{(\lambda-1)}(I+x\tilde{M}).$
Each entry of $(s+1)C_{s}^{(\lambda-1)}(I+x\tilde{M})u(t)$ is a non-negative finite linear combination of $[u]_{j}$’s by the RKL1 monotonicity; by the inductive hypothesis, each entry of $C_{s-1}^{(\lambda)}(I+x\tilde{M})u(t)$ is also a non-negative finite linear combination of $[u]_{j}$’s. Finally, for $0\leq x\leq\frac{1}{2}$, each entry of $(I+x\tilde{M})v$ is again a non-negative finite linear combination of $[v]_{j}$’s for any $v\in\mathbb{R}^{\mathbb{Z}}$; therefore, each entry of $(I+x\tilde{M})C_{s-1}^{(\lambda)}(I+x\tilde{M})u(t)$ is a non-negative finite linear combination of $[u]_{j}$’s. Taking the summation of two terms, each entry of $C_{s}^{(\lambda)}(I+x\tilde{M})u$ is a non-negative finite linear combination of $[u]_{j}$’s; hence RKG1 is monotone for the desired CFL condition.
@@
The $s$-stage RKG2 method 14 applied to constant coefficient heat equation under periodic or homogeneous Neumann boundary condition is monotone if it satisfies its stability constraint
@@
(a) The $s$-stage RKG1 method 13 applied to constant coefficient semilinear heat equation with Strang splitting is monotone if it satisfies its stability constraint
@@
(b) The $s$-stage RKG2 method 14 applied to constant coefficient semilinear heat equation with Strang splitting is monotone if it satisfies its stability constraint
@@ Appendix B Numerical Discretization
(26) $(\mathbf{M}r)_{m}^{n}$ $:=\frac{1}{2r_{m}^{n}dx}\Biggl{[}\frac{r_{m+1}^{n}}{\sqrt{1+(\delta r_{m+1}^{n})^{2}}}\Biggl{(}\frac{\tilde{\mathcal{H}}_{m+2}^{n}-\tilde{\mathcal{H}}_{m}^{n}}{2dx}\Biggl{)}$
@@
(27) ~$\tilde{\mathcal{H}}_{m}^{n}$ $:=\frac{1}{r_{m}^{n}\sqrt{1+(\delta r_{m}^{n})^{2}}}-\frac{\delta^{2}r_{m}^{n}}{(1+(\delta r_{m}^{n})^{2})^{3/2}}$
$\delta r_{m}^{n}$ $:=\frac{r_{m+1}^{n}-r_{m-1}^{n}}{2dx}\quad\delta^{2}r_{m}^{n}:=\frac{r_{m+1}^{n}-2r_{m}^{n}+r_{m-1}^{n}}{dx^{2}}$
@@ Appendix C Numerical Monotonicity Example
where $dt$ represents the timestep and $dx$ represents the spatial grid size. Then in this case, $j$ enumerates from $i-1$ to $i+1$ for a fixed grid index $i$; more explicitly,
@@
Throughout our proof, we focus on the problem 19. In that case, for a $s$-stage stabilized Runge–Kutta method, $j$ enumerates from $j-s$ to $j+s$.
@@
Since $14x<$ 1, using Theorem 5.7 with $a=\frac{j+s+1}{2},b=\frac{j-s}{2}$, we know that

### v2

@@ 5.2.1 Combinatorics Formulation
when applied to problem 18 if and only if
@@
Plugging into the RKL1 method 10, we have that
@@
where $M^{k}u^{t}]_{i}$ denotes the $i$-th element of the vector $M^{k}u^{t}$. Using induction, it can be shown that
$M^{k}u^{t}]_{i}=\frac{1}{dx^{2k}}\sum_{j=0}^{2k}\binom{2k}{j}(-1)^{j}u_{i+j-k}^{t}=\frac{1}{dx^{2k}}\sum_{j=-k}^{k}\binom{2k}{j+k}(-1)^{j+k}u_{i+j}^{t}.$
@@ 5.2.2 Estimation of the Alternating Sum and Conclusion
In particular, it gives us a form of generalized hypergeometric functions that is always non-negative. Our goal is to write the alternating sum 19 as a generalized hypergeometric function, with the hope that the coefficients are compatible with Clausen’s formula and directly apply the theorem to conclude. Fortunately, this process indeed gives us the desired result using the calculations in the following lemma.
@@
(a) The $s$-stage RKL1 method 10 applied to constant coefficient heat equation under periodic or homogeneous Neumann boundary condition is monotone if and only if
@@
(b) The $s$-stage RKL2 method 11 applied to constant coefficient heat equation under periodic or homogeneous Neumann boundary condition is monotone if and only if
@@
(a) The $s$-stage RKL1 method 10 applied to constant coefficient semilinear heat equation with Strang splitting is monotone if and only if it satisfies the stability constraint
@@
(b) The $s$-stage RKL2 method 11 applied to constant coefficient semilinear heat equation with Strang splitting is monotone if and only if it satisfies the stability constraint
@@ 5.3 Runge–Kutta–Gegenbauer Methods Monotonicity
The $s$-stage RKG1 method 12 applied to constant coefficient heat equation under periodic or homogeneous Neumann boundary condition is monotone if it satisfies its stability constraint
@@
Also notice that $C_{s}^{(1/2)}(z)=P_{s}(z)$ is exactly the $s$-th degree Legendre polynomial. The $s$-stage RKL1 monotonicity is equivalent to the statement that $\Big[P_{s}\Big(I+x\tilde{M}\Big)u(t)\Big]_{i}$ is a non-negative finite linear combination of $u]_{j}$’s, where $\tilde{M}u]_{i}:=[u]_{i+1}-2[u]_{i}+[u]_{i-1}$, under the condition $0\leq x\leq\frac{1}{2}$.
@@
[…]
Each entry of $(s+1)C_{s}^{(\lambda-1)}(I+x\tilde{M})u(t)$ is a non-negative finite linear combination of $u]_{j}$’s by the RKL1 monotonicity; by the inductive hypothesis, each entry of $C_{s-1}^{(\lambda)}(I+x\tilde{M})u(t)$ is also a non-negative finite linear combination of $u]_{j}$’s. Finally, for $0\leq x\leq\frac{1}{2}$, each entry of $(I+x\tilde{M})v$ is again a non-negative finite linear combination of $v]_{j}$’s for any $v\in\mathbb{R}^{\mathbb{Z}}$; therefore, each entry of $(I+x\tilde{M})C_{s-1}^{(\lambda)}(I+x\tilde{M})u(t)$ is a non-negative finite linear combination of $u]_{j}$’s. Taking the summation of two terms, each entry of $C_{s}^{(\lambda)}(I+x\tilde{M})u$ is a non-negative finite linear combination of $u]_{j}$’s; hence RKG1 is monotone for the desired CFL condition.
@@
The $s$-stage RKG2 method 13 applied to constant coefficient heat equation under periodic or homogeneous Neumann boundary condition is monotone if it satisfies its stability constraint
@@
(a) The $s$-stage RKG1 method 12 applied to constant coefficient semilinear heat equation with Strang splitting is monotone if it satisfies its stability constraint
@@
(b) The $s$-stage RKG2 method 13 applied to constant coefficient semilinear heat equation with Strang splitting is monotone if it satisfies its stability constraint
@@ Appendix B Numerical Discretization
(25) $(\mathbf{M}r)_{m}^{n}$ $=\frac{1}{2r_{m}^{n}dx}\Biggl[\frac{r_{m+1}^{n}}{\sqrt{1+(\delta r_{m+1}^{n})^{2}}}\Biggl(\frac{\tilde{\mathcal{H}}_{m+2}^{n}-\tilde{\mathcal{H}}_{m}^{n}}{2dx}\Biggl)$
@@
(26) $\tilde{\mathcal{H}}_{m}^{n}$ $=\frac{1}{r_{m}^{n}\sqrt{1+(\delta r_{m}^{n})^{2}}}-\frac{\delta^{2}r_{m}^{n}}{(1+(\delta r_{m}^{n})^{2})^{3/2}}$
$\delta r_{m}^{n}$ $=\frac{r_{m+1}^{n}-r_{m-1}^{n}}{2dx}\quad\delta^{2}r_{m}^{n}=\frac{r_{m+1}^{n}-2r_{m}^{n}+r_{m-1}^{n}}{dx^{2}}$
@@ Appendix C Numerical Monotonicity Example
where $dt$ represents the timestep and $dx$ represents the spatial grid size. Then in this case, $j$ enumerates from i1i-1 to $i+1$ for a fixed grid index $i$; more explicitly,
@@
Throughout our proof, we focus on the problem 18. In that case, for a $s$-stage stabilized Runge–Kutta method, $j$ enumerates from jsj-s to $j+s$.
@@
Since $4x<1$, using Theorem 5.7 with $a=\frac{j+s+1}{2},b=\frac{j-s}{2}$, we know that